Diffusion Morning Report 09/25/08

Ted Berg tberg at stanford.edu
Thu Sep 25 06:16:46 PDT 2008


Good Morning All,
    Looking OK this morning . LTO tube change in progress. Ray is out 
today and tomorrow. Mike is helping out-Thanks. TEOS2v had leak alarms. 
Nitride had MFS alarms. Thermco poly2 has non uniform dep. Ebeam 
replaced tip. No other problems at this time.

  	  	  	  	  	 
  	Date: 	9/25/08 	THICKNESS-Change at 	  	 
Tube Name/Number 	Status 	  	  	  	 
Tylan 1 	up 	  	  	  	 
Tylan 2 	up 	  	  	  	 
Tylan 3 	up 	  	  	  	 
Tylan 4 	up 	  	  	  	 
Tylan 5-BBr3 	up 	  	  	  	 
Tylan 6-POCl3 	up 	  	  	  	 
Tylan 7-TEOS2 	yellow 	TEOS leak alarm 	18.16/30 	  	 
Tylan 8-SiGe 	up 	  	25.93/30 	  	 
Tylan 9-Poly 	up 	  	1.4/30 	  	 
Tylan 10-Nitride 	yellow 	mfs alarms 	2.39/30 	  	 
Tylan 11-LTO 	down 	contamination-tube change 	0/30 	  	 
Tylan 12-FGA 	up 	  	  	  	 
Thermco 1 	up 	  	  	  	 
Thermco 2 	up 	  	  	  	 
Thermco 3 	future 	  	  	  	 
Thermco4 	future 	  	  	  	 
Thermco Poly1 	up 	  	  	  	 
Thermco Poly2 	yellow 	non uniformity 	  	  	 
Thermco Nitride1 	up 	  	  	  	 
Thermco Nitride2 	future 	  	  	  	 
TEL Spa 	up 	  	  	  	 
Burn Box 	up 	
	   	  	 
Ebeam 	down 	Replaced tip 	  	  	 
ZYGO 	up 	  	  	  	 
FGA2 	up 	  	  	  	 
SiH4 for furnaces 	  	B on line 650 psi 	 changeover working well  	  	 


-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20080925/58691005/attachment.html>


More information about the diffusion mailing list